OpenStudy (anonymous):

Find the domain of the logarithmic function f(x) = log4 (x – 3).

OpenStudy (anonymous):

A.Any real numbers B. Positive real numbers C.Negative real numbers D.x > 3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Can you take the log of any number?

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

No. Can you take the log of any number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait no no u cant

OpenStudy (anonymous):

That's incorrect. For example, you cannot take the log of -1; because 10 to no power will give you -1. (I was using base 10..but makes no difference what base you are using).

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

So the underlining question is...what kind of numbers can you take the log of?

OpenStudy (anonymous):

all reall numbers

OpenStudy (anonymous):

In your probelm, x cannot be -7, because x-3 will be -7-3 = -10, and you can't take the log of -10.

OpenStudy (anonymous):

You cannot take the log of all real numbers. We just discussed that you cannot take the log of -1. Back to our initial question...think as to what numbers can you take the log of?

OpenStudy (anonymous):

positive real numbers?

OpenStudy (anonymous):

exactly! You can't even take the log of 0....because 10 to no power gives you 0. So since you can only take the log of a positive real number, then in your function, we better make sure that x - 3 > 0.

OpenStudy (anonymous):

So what will be domain of y = log (x-3)?

OpenStudy (anonymous):

what are all the numbers x in the domain for y = log(x - 3)?

OpenStudy (anonymous):

numbers 1+ or all positive real numbers

OpenStudy (anonymous):

that is true for y = log x we are looking at the function y = log (x - 3). Since we can only take the log of a positive number, then we must make sure that x - 3 > 0!! Agree?

OpenStudy (anonymous):

yeah my bad i didnt see the question right i cant put 1 or 2

OpenStudy (anonymous):

so whats the right answer to this question?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

x>3

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

wow your really smart

OpenStudy (anonymous):

Thanks for help

OpenStudy (anonymous):

It has nothing to do with smart. It has to do with understanding the log function...that we can only take the log of a number greater than 0 (regardless of the base).

OpenStudy (anonymous):

And to understand why that is true; not memorize.

OpenStudy (anonymous):

Look at the graph of y = log (x-3).

OpenStudy (anonymous):

what about a number greater then or equal to 0 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

you can see from the graph, that x exists for x>3, only.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK. Good luck!

OpenStudy (anonymous):

i got a other one i did y=1/x-2 and i put all real numbers exept 2 for a awnser

OpenStudy (anonymous):

Is the entire x-2 in the denominator? y = 1/(x-2)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you are correct, if asked for the domain. x van be any real number other than 2, as the function is undefined at x = 2.

OpenStudy (anonymous):



OpenStudy (anonymous):

So you are correct for that problem.

OpenStudy (anonymous):

thanks:D

OpenStudy (anonymous):

welcome.